A man is bowling and the probability of him knocking all the pins is 0.3. If he has 10 shots, what is the probability that hell knock all the pins down less than 3 times?


1
Expert's answer
2022-05-30T16:57:39-0400

Let "X=" the number of times a man knocks all the pins down: "X\sim Bin(n, p)."

Given "n=10, p=0.3, q=1-p=0.7."


"P(X<3)=P(X=0)+P(X=1)+P(X=2)"

"=\dbinom{10}{0}(0.3)^{0}(0.7)^{10-0}+\dbinom{10}{1}(0.3)^{1}(0.7)^{10-1}"

"+\dbinom{10}{2}(0.3)^{2}(0.7)^{10-2}=0.3827827864"